OpenStudy (anonymous):

help please 2x+5y=-13 , 3x-4y=-8

OpenStudy (anonymous):

5y= -13 -2x y= -13/5 -2x/5 3x - 4(-13/5 -2x/5) = -8 3x + 10.4 + 1.6x =-8 4.6x = -18.4 x = -4 2(-4) + 5y = -13 -8 +5y = -13 5y = -5 y = -1